Write the equation of a line passing through point (0, -4) with a slope of (-3/4).
Answer by mananth(16946)

the general equation in point slope form is
(y-y1)=m(x-x1)
whwre (x1,y1) is the point and m is the slope
(0,-4) & (-3/4)=m
(y-(-4)=(-3/4)(x-0)
y+4 = -3/4 x
y= (-3/4)x -4